Suppose that the cost function of a firm is C(q)=4q. Suppose that this is the only firm in the market, and demand is Q(p)=10-p. What is the Producer Surplus in a competitive equilibrium in this economy if the government imposes an advalorem tax of 20% (this means that if the firm sells at price p, the government receives 0.2p per unit sold)?

6

12

1

0

2
Answer #1

New demand Curve

Q = 10-1.2P

1.2P = 10-Q

P = (10/1.2) - Q/1.2

MR = (10/1.2) - 2*(Q/1.2)

MR = (10/1.2) - Q/.6

New eqm,

MR = MC, as single firm case

(10/1.2) - Q/.6 = 4

Q = 4.33*.6 = 2.6

New market price = 10-2.6 = 7.4

PS = π = (7.4-4)*2.6

=